OpenStudy (darkbeautystar):

2 Swimming Pools have the same area. One is circular and has a diameter of 12 feet. The other is a parallelogram with a height of 3 feet. What is the length of pool that is shaped like a parallelogram?

OpenStudy (phi):

do you know how to find the area of a circle? what is the formula ?

OpenStudy (darkbeautystar):

I already know the area of the circle is 113.1

OpenStudy (phi):

ok. Just to be sure, do they give the volume of either pool ?

OpenStudy (darkbeautystar):

No

OpenStudy (phi):

any more info? The reason I ask, is they say the "height" of the second pool. unless that is the width, we can't do the problem

OpenStudy (darkbeautystar):

So here is the info I have: Both pools have the same area The circle pool has a diameter of 12 feet The other one is a parallelogram with a height of 3 feet

OpenStudy (darkbeautystar):

And it wants to know the length of the parallelogram pool

OpenStudy (darkbeautystar):

I'm pretty sure that it is 37.7

OpenStudy (phi):

If that is all the info there is, this is a terrible question. if we have to assume they made a mistake and mean the width is 3 then we do 3*x= 113.1 x= 113.1/3 = 37.7 (or more exactly 12pi)

OpenStudy (misssmartiez):

Sorry I was late, I was helping someone, looks likes you have it covered though!